log base 2(x-1)+log base 6(x-1)=2 solve for x
Found 2 solutions by lynnlo, MathTherapy:
Answer by lynnlo(4176)

x=1+2/log(12)

Answer by MathTherapy(10809)

log base 2(x-1)+log base 6(x-1)=2 solve for x

The other person's answer: x=1+2/log(12), is WRONG!!

Correct answer: 

Send comments and "thank-yous" to "D" at MathMadEzy@aol.com